OpenStudy (anonymous):

An actor invests some money at 8%, and 50,000 more than three times the amount at 9% . The total annual interest earned from the investment is 37,750. How much did he invest at each amount? He invested$____at 8% and$____at 9%

OpenStudy (anonymous):

a= amt @ 8% b= amt @ 9% b= 3a+50000 .08a+.09b=37750 .08a+.09(3a+50000)=37750 .08a+.27a+4950=37750 .35a=32800 a=93714.29 b=331142.86

OpenStudy (anonymous):

always double check with that many numbers